answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Can you have a period 4 days after conception?

If you truly conceived, no. Women do not have regular periods after becoming pregnant. What you experienced was probably implantation bleeding. This bleeding is caused by the fertilized egg attaching itself to the uterine wall.

How do you solve for g in a simple pendulum problem?

For a simple pendulum, with very small, technically infinitesimal, swings, the period T is approximately 2 pisquare root (Length / gravity). From there, it is simple algebra.

T2 = 4 pi2 length / gravity

gravity = 4 pi2 length / T2

For more information, including how to compensate for circular error introduced by non-infinitesimal swings, please see the Related Link below.

What is hyper-geometic probability?

The hyper-geometric distribution is a discrete probability distribution which is similar (in some respects) to the binomial distribution. Suppose you have a population of N which contains R successes. The Binomial describes the probability of r successes in n draws out on N with replacement.

However, in many situations the draw is not replaced. In this case you get the hyper-geometric distribution.

The function is given by:

Prob(r successes in n draws out of N) = RCr/[N-RCn-r * NCn]

With the binomial distribution the probability of success remains constant (=R/N) throughout. With the hypergeometric, the numerator for success reduces by one after each successful outcome whereas the denominator reduces by one whatever the outcome.

What is the probability of getting exactly 3 heads if a fair coin is tossed 8 times?

The probability to get heads once is 1/2 as the coin is fair

The probability to get heads twice is 1/2x1/2

The probability to get heads three times is 1/2x1/2x1/2

The probability to get tails once is 1/2

The probability to get tails 5 times is (1/2)5

So the probability to get 3 heads when the coin is tossed 8 times is (1/2)3(1/2)5=(1/2)8 = 1/256

If you read carefully you'll understand that 3 heads and 5 tails has the same probability than any other outcome = 1/256

As the coin is fair, each side has the same probability to appear

So the probability to get 3 heads and 5 tails is the same as getting for instance 8 heads or 8 tails or 1 tails and 7 heads, and so on

What is the probability that you will pull out a heart in a deck of cards?

There are 13 hearts in a 52 card deck. The probability of selecting a heart out of the deck is 13/52 or 1/4 or 0.25.

When a family of seven boys and no girls is expecting an eighth child what is the probability of the next child being a girl?

If the gender of a child were an independent variable then the genders of the existing children would be irrelevant and so the probability of the next child being a girl would be approximately 1/2.

It would be approximately 1/2 because the overall proportion is not exactly half. However, and more important, is the fact that the gender of a child is affected by the parents' genes and so is not independent of the gender of previous children.

What is the probability of drawing below a five in a deck of cards?

The answer depends on whether the card(s) are drawn at random, whether it is a standard deck (or a pinochle deck), how many cards are drawn, if more than one then whether they are replaced before the next one is drawn.

If 40 cards are drawn, without replacement, from a standard deck of cards, the probability of a card below 5 is 1: that is, a certainty.

For a single random draw from a standard deck, the probability is 3/13 (Aces high) or 4/13 (Aces low).

How is expected rate of return calculated from average rate of return on investment and standard deviation?

The expected rate of return is simply the average rate of return. The standard deviation does not directly affect the expected rate of return, only the reliability of that estimate.

What is the probability of UFO's existing?

100%. UFOs can be any unidentified flying objects. They don't necessarily have to be driven by aliens.

What is the probability of a simultaneous death of spouses who are both over age 65?

Minuscule.

Deaths are usually separated by seconds, minutes or longer time periods.

What is certain is that they will eventually die.

What are the odds of having twins with no family history?

Every one who has ever been born has a family history, no matter what their circumstance.

What are the probable sources of smoke?

sources of smoke include emissions from vehicles and power stations

What is the probability of a coin landing on heads when it is flipped 100 times?

The probability of landing on heads at least once is

1 - (1/2)100 = 1 - 7.9*10-31 which is extremely close to 1: that is, the event is virtually a certainty.

How do you calculate the probability of an event occurring at least once over three attempts if the likelihood of success increases on each attempt?

Not knowing exactly what the 'increase' is, it is hard to answer this. Here is what happens if the event probability is the same each time (at the bottom is one example where probability increases):

Let p = probability that the event will occur on a trial (attempt).

Let q = probability against the event occurring on a trial = 1 - p. [100% - probability for]

Probability that the event will not occur on n trials = qn, so the probability that the event will occur at least once on n trials is 1 - qn = 1 - (1-p)n.

Example: Say you roll a 6-sided die twice and want the probability that you'll get a 1 at least one of the throws.

p = 1/6, n = 2 : Probability = 1 - (1 - 1/6)² = 30.56%

Let's look at another way to figure this. Rolling a die twice, it's easy enough to make a 2-dimensional table of the results. There are 36 possibilities, and 11 of them have a 1 occurring. So the probability is 11/36 = 30.56%

If you were to do 3 attempts, you would have a 3-dimensional table with 216 possibilities, and 91 of them have a 1 occurring at least once.

91/216 = 42.13%, which is the same using the formula: 1 - (1 - 1/6)³ = 42.13%

So for the problem where the probability increases each time. As an example, say you're playing BINGO with 75 balls. Say you're playing a blackout game and 50 balls have already been drawn, and you've filled your block except for a single square (say it's number 1). So there are 25 balls remaining. It will be similar to above, except you cannot just square or cube the probabilities because they change.

You still want to calculate the probabilities against, then multiply them together then subtract from 100%. Since this is an example, you're playing BINGO and you look around at several neighbors and see that everybody around you has three spaces to fill, so you want to hit #1 in the next three balls.

Assign the probabilities for picking #1 on each trial as p1, p2, p3.

So p1 = 1/25. If you don't draw #1 the first time, then there are only 24 balls left, so p2 = 1/24, and p3 = 1/23. We need to multiply the q's (probabilities against) then subtract that product from 100%.

So the probability of drawing #1 on any of the next 3 trials, starting with 25 balls is: 1 - ((1 - 1/25)(1 - 1/24)(1 - 1/23))= 1 - ((24/25)*(23/24)*(22/23)) = 12%. Notice the way I wrote it, the fractions cancel. Here's the interesting thing:

The probability of getting #1 on the first try is 1/25 = 4%. The first or second try is 1 - ((24/25)*(23/24)) = 2/25 = 8%. Each trial will become an increased number over 25: 3/25 = 12%, 4/25 = 16%, etc. I'm not sure how you would generalize the formula though.

What is the probability of rolling a 6 and a 2.?

The key word here is and. So both events have to happen. If you're rolling two dice, then die A can have a 6, and die B can have a 2, OR die A can have a 2 and die B has a 6. These are the 2 'winning' outcomes from a total of 36 possible outcomes with two dice. 2 out of 36 = 1/18 = 5.56%

2 What probably happen if an astronauts suit spacesuit was black?

it would over heat. Black objects typically have a higher solar heat gain coefficient than lighter colors.

When do you use theoretical probability?

When you have a model that describes the required situation well enough so that you can use scientific laws to calculate the probabilities of the outcomes of the experiment.

What is an event with a probability of zero?

mars having zebras as TV hosts for their segment on Earth

Trending Questions
What is the probability that a random selected poker hand contains exactly 3 aces given that it contains at least 2 aces? How many ways can Cindy arrrange 42 books in different shelf if each shelf has to be the same? When did Odd Isøy die? What is the probability of choosing a marble that is not blue in problem 2? Is the chance of getting a 2 or 4 in rolling a die? What are the odds on a single roll of a pair of dice against rolling a sum of 6? Is Probability the chance that something will or will not happen? What is the probability of rolling a 7 with one dice? If person A can hit a target 6 times in 7 shots and B can hit it 4 times in 5 and C can hit it 3 times in 4 then what is the probability the target will be hit exactly twice if A B and C take one shot? A typical expectation in actual practice is that estimates should have a probability of being met of the time? What does without replacement mean in probability? The probability of 32 out of 80? What is the possibility of rolling a three if you roll 4 dices at the same time? What is the probability of drawing a king and a face card? How many possible outcomes when a standard six sided die is rolled? What is the probability of getting an even number when throwing a die? What process is shown occurring at C and what does it occur? What is the luckiest number if you roll 3 dice? When two dice are rolled what is the probability of getting a sum greater than or equal to 9? What is a requirement for use of the hypergeometric distribution Only 2 possible outcomes Trials are independent Probability of a success is greater than 1.0 Sampling with replacemen?